# Cheng Lin4, Shiqing Xin5, Yuan Liu3†, Wenping Wang6, Taku Komura1

### v1

@@ Abstract
Scaling artist-designed meshes to high triangle numbers remains challenging for autoregressive generative models. Existing transformer-based methods suffer from long-sequence bottlenecks and limited quantization resolution, primarily due to the large number of tokens required and constrained quantization granularity. These issues prevent faithful reproduction of fine geometric details and structured density patterns. We introduce MeshMosaic, a novel local-to-global framework for artist mesh generation that scales to over 100K triangles—substantially surpassing prior methods, which typically handle only around 8K faces. MeshMosaic first segments shapes into patches, generating each patch autoregressively and leveraging shared boundary conditions to promote coherence, symmetry, and seamless connectivity between neighboring regions. This strategy enhances scalability to high-resolution meshes by quantizing patches individually, resulting in more symmetrical and organized mesh density and structure. Extensive experiments across multiple public datasets demonstrate that MeshMosaic significantly outperforms state-of-the-art methods in both geometric fidelity and user preference, supporting superior detail representation and practical mesh generation for real-world applications. Our project is available at: https://xrvitd.github.io/MeshMosaic/index.html.
[…]
@@ 1 Introduction
Artist-designed triangular meshes are a cornerstone in film, gaming, AR/VR, and industrial design. High-quality artist meshes are central to computer graphics and 3D vision, distinguished by their stylized topology, directional flows, uneven triangle densities, sharp edges, and symmetry. Recent advances in 3D generation (Xiang et al., 2025; Li et al., 2024a; Liu et al., 2023; Long et al., 2024) and reconstruction (Wang et al., 2023; Huang et al., 2024; Xu et al., 2023; Hou et al., 2022) highlight the limitations of classical meshing methods like Marching Cubes (Lorensen & Cline, 1998), which rely on uniform grids and produce redundant triangles, struggling with sharp features. Traditional meshing either yields uniform (Liu et al., 2009; Xu et al., 2024; Wang et al., 2025b; Dong et al., 2025b; a) or oversimplified (Chen et al., 2023; Garland & Heckbert, 1997) results; while anisotropic techniques (Zhong et al., 2014) better align to curvature, they still fall short in capturing the varying density and structure of artist meshes.
Figure 2: Mosaic Art (McPhee, 2025).
@@
The rise of large language models (LLMs) (Zhao et al., 2023a) has inspired GPT-like architectures for mesh generation, such as MeshGPT (Siddiqui et al., 2024) and its successors (Chen et al., 2024a; c; Zhao et al., 2025; Hao et al., 2024; Tang et al., 2024b). Despite progress, these approaches struggle to scale up due to prohibitively long token sequences and limited quantization resolution, making it difficult to generate high-triangle meshes with fine detail. However, in practice, artist-designed meshes often require significantly higher resolutions to achieve the visual fidelity demanded in modern games and films. For example, production-quality character models or hero assets frequently contain upwards of 100K faces, far exceeding the capacities handled by current generative methods (Zhao et al., 2025; Weng et al., 2025; Lionar et al., 2025). This substantial gap underscores the need for methods capable of generating high-triangle meshes that preserve the intricate details and structural coherence.
@@
Inspired by the compositional principles of classical mosaic art (Fig. 2), we propose MeshMosaic, a novel local-to-global framework for scalable artist mesh generation. Mosaic artworks achieve global complexity and coherence by assembling intricate local tiles; in a similar spirit, MeshMosaic constructs a complete mesh by stitching together multiple locally generated patches. Unlike previous methods that attempt to model the entire mesh sequence, our framework divides the mesh into semantically meaningful patches, each autoregressively generated from a full-size point cloud with full-resolution quantization. To enable a compact yet faithful geometric representation, we employ shared boundary conditions and semantic segmentation, which address challenges related to boundary alignment and asymmetry. This patch-based strategy not only sidesteps the long-sequence bottleneck, but also effectively captures fine-grained geometric structures and global coherence, allowing for high-detail modeling within each patch while maintaining consistency across the entire mesh.
@@
Experiments on multiple datasets show that MeshMosaic establishes new milestones in geometric fidelity and detail, also strongly preferred in user studies for artistry. Our approach supports stable generation of high-resolution meshes with over 100K triangles (see Fig. 1) and faithfully reproduces fine detail via per-patch quantization. See Fig. 3 for a gallery of our results.
@@ 2 Releated Works
Remarkable advances have been made in 3D shape generation, particularly with the adoption of signed distance field (SDF) representations, which offer notable accuracy and flexibility for modeling complex shapes. Despite such progress, these SDF-based methods often depend on the Marching Cubes algorithm (Lorensen & Cline, 1998) for mesh extraction, which can result in redundant triangles and consequently large file sizes—posing limitations for scalable deployment and real-time applications.
@@
Figure 3: Gallery of our artist mesh generation results.
For instance, Wonder3D (Long et al., 2024) introduces a cross-domain diffusion framework for generating high-quality, multi-view textured 3D meshes from single images, achieving improved consistency and visual fidelity over previous approaches. CLAY (Zhang et al., 2024) expands the scope with a large-scale generative model that transforms text, images, and 3D-aware inputs into intricate geometry and material compositions, making robust 3D asset creation accessible to broad user bases. TRELLIS (Xiang et al., 2025) leverages structured occupancy fields to guide the formation of salient shape features, supporting high-precision modeling conditioned on text or image prompts. Hunyuan3D-2.5 (Lai et al., 2025) proposes a two-stage diffusion pipeline for crafting high-fidelity assets, combining powerful generative models with physically-based rendering for enhanced realism in both shape and texture. CraftsMan3D (Li et al., 2024b), evolves toward interactive 3D design by developing a native diffusion-based framework capable of producing meshes with regular topology and fine surface detail, while supporting user-driven refinements.
2.2 Artist Mesh Generation
The quest for artist-quality mesh generation has inspired a new wave of models that focus on efficient topology and expressive geometry. MeshGPT (Siddiqui et al., 2024) pioneers autoregressive mesh synthesis through sequence-based modeling, employing quantized latent embeddings and transformer architectures to predict efficient triangulation and structural patterns reminiscent of hand-crafted meshes. Building on this idea, MeshAnything (Chen et al., 2024b) and MeshAnythingV2 (Chen et al., 2024c) offer advanced mesh generation using adjacent mesh tokenization, reducing token sequence lengths and enabling more complex, artist-grade meshes, with MeshAnythingV2 doubling the operational face limit.
MeshXL (Chen et al., 2024a) introduces the Neural Coordinate Field, which fuses explicit coordinate representation with implicit neural embeddings for more scalable, high-fidelity mesh modeling. EdgeRunner (Tang et al., 2024b) addresses past limitations of autoregressive mesh approaches by presenting an improved tokenization algorithm and compressing variable-length meshes into fixed-size latent vectors, yielding more diverse, generalizable, and higher-quality outputs. Meshtron (Hao et al., 2024) leverages a novel hourglass neural architecture with sliding window inference and robust sampling, achieving new levels of scalability and fidelity.
In addition, TreeMeshGPT (Lionar et al., 2025) introduces a tree sequencing method for triangle adjacency, dynamically growing mesh structures during autoregressive generation for improved training and mesh quality. iFlame (Wang et al., 2025a) balances efficiency and generative power by combining linear and full attention within an hourglass framework, augmented by caching for fast inference and training. Nautilus (Wang et al., 2025c) explores locality-aware autoencoding by leveraging manifold mesh properties, novel tokenization, and dual-stream conditioning, significantly enhancing scalability and structural consistency.
Compression-oriented approaches such as Blocked and Patchified Tokenization (BPT) (Weng et al., 2025) further reduce token sequence length, allowing detailed mesh synthesis with more faces. Building on BPT, DeepMesh (Zhao et al., 2025) integrates reinforcement learning for human preference alignment, supporting the generation of intricately detailed meshes with precise topology. In addition to autoregressive-based approaches, methods such as PolyDiff (Alliegro et al., 2023) and PDT (Wang et al., 2025b) directly employ diffusion models to generate structured triangles or points from Gaussian noise.
@@ 2.3 Part-based Shape Generation
Part-based shape generation rests on the principle that decomposing objects into semantic parts furnishes rich priors for structure-aware reconstruction and controllable synthesis. Universal segmentation techniques have scaled part discovery across a wider range of data. Segment Any Mesh (Tang et al., 2024a) generalizes promptable segmentation to 3D meshes, supporting flexible and category-agnostic part extraction crucial for interactive and generative workflows. SAM3D (Yang et al., 2023) adapts this paradigm to large-scale 3D scenes, enabling multi-granular, prompt-driven segmentation. By distilling knowledge from SAM’s multi-view segmentation results, SAMPart3D (Yang et al., 2024) further specializes in part segmentation for individual objects. More recently, PartField (Liu et al., 2025) represents shapes as continuous feature fields and trains a transformer-based feed-forward network with an ambiguity-agnostic contrastive loss, achieving efficient and high-quality open-world part segmentation. PartCrafter (Lin et al., 2025) jointly creates multiple semantically distinct parts from a single image, enabling end-to-end part-aware 3D mesh synthesis with global coherence and fine-grained detail.
Human modelers typically create models based on their understanding of component-based structures (Lin et al., 2020), and thus part-based generation is a problem of significant importance. Part123 (Liu et al., 2024) illustrates this by reconstructing 3D shapes from single images while predicting semantic parts and their spatial arrangement. ComboStoc (Xu et al., 2024) introduces combinatorial stochasticity into diffusion by jointly sampling discrete structural decisions (such as part templates or multiplicity) with continuous geometry. These segmentation frameworks underpin part-based shape generation by providing scalable, promptable part vocabularies and supporting interactive conditioning and evaluation at the part level.
[…]
Figure 4: The pipeline of MeshMosaic. During inference, our method first applies PartField (Liu et al., 2025) to obtain semantic segmentation of the input shape. The input point cloud is then sampled according to the segmented patches and the original shape. Finally, our approach produces a clean, highly detailed mesh by assembling the generated patches.
@@ 3.2 Local-to-Global Mesh Generation
We use PartField (Liu et al., 2025) for semantic segmentation at inference time (see Fig. 4), which embeds semantic structure and produces well-aligned boundaries, often guided by curvature flow to enhance realism and make future edits easier.
Sorting patches.
Then, we will generate the whole mesh in a part-by-part manner, which requires us to determine a generation order. Thus, the patch generation is carried out in breadth-first search (BFS) order, beginning from the spatially lowest patch and then proceeding to adjacent patches. Sequential generation with the autoregressive model ultimately yields the final mesh assembly. Fig. 5 shows a 2D illustration with eight patches, where the black line demonstrates the mesh surface. Then, for each patch, we adopt the following structure to generate the triangle meshes.
@@
Figure 6: The workflow of MeshMosaic for generating a single patch. Both global and local point cloud features are extracted by a locked Michelangelo (Zhao et al., 2023b) encoder. For each patch, the nearest boundary mesh is identified, tokenized, and concatenated before the target mesh token sequence. The GRU network encodes boundary tokens, which are then combined with global and local features and fed into an autoregressive hourglass transformer for mesh generation.
@@ 3.3 Generating single Patch
When generating triangles for a specific patch, we will use the triangles from existing generated patches as boundary conditions. This essentially enables the smooth connection between different patches. Specifically, we introduce an efficient boundary conditioning mechanism: the token sequence from earlier patches is fed as context to subsequent patches. To avoid inefficiency and information dilution from excessively long token sequences, for each patch, we select only 512 spatially nearest triangles from prior patches. These are tokenized, passed through a Gate Recurrent Unit (Cho et al., 2014) (GRU) network, and the resulting embedding conditions the transformer network (see Fig. 6, blue arrows). Fig. 7 shows an example of the plane shape. The whole shape was segmented into three patches. Each patch gets the boundary information from the previous patches, following the BFS order. For the very first patch, where no previous boundaries exist, we supply a placeholder token sequence consisting entirely of terminator tokens to the GRU network, establishing a neutral starting context for the generation process.
@@
Local-to-Global Conditioning.
While boundary conditioning enforces local continuity, we enhance global coherence via local-to-global point cloud features (middle of Fig. 6). During training and inference, our autoregressive model is conditioned on representations from both the current patch point clouds and the full shape point clouds. Both point clouds are encoded with a frozen Michelangelo (Zhao et al., 2023b) encoder. The extracted global and local features are concatenated with GRU boundary features and provided to the transformer as final condition input.
Local Quantization.
Given both the boundary conditions and local-to-global information, we generate the triangles locally using a local quantization. Unlike previous approaches, such as DeepMesh (Zhao et al., 2025), which apply a uniform quantization of $512^{3}$ resolution to the entire mesh, our method independently scales each patch to $[0,1]$ and quantizes it at $512^{3}$ resolution. This local quantization approach enables a higher effective merged resolution, allowing for the preservation and recovery of richer geometric details. As shown in Fig. 4 and Fig. 6, our pipeline first segments the shape into patches, with each patch quantized independently to $512^{3}$ resolution and provided with 16,384 sampled points as input. In contrast to baseline methods, which quantize the full shape and use only a single set of 16,384 point cloud samples, our framework assembles meshes from individually quantized patches, each paired with its respective sampled points. This strategy offers the dual benefits of higher overall shape resolution and a greater abundance of conditional information for the network to leverage during generation.
Gluing local patches.
It should be noted that local quantization may introduce minor positional displacements for each patch, which can lead to discontinuities along patch seams if not properly addressed. To ensure seamless integration, we compute the displacement between the position of boundary condition faces referenced by the current patch and their corresponding original quantized positions in the previously assembled patches. The entire current patch is then translated according to this computed displacement, aligning it precisely with previously assembled patches. This compensatory adjustment guarantees smooth boundaries and continuity across the entire mesh, enabling high-fidelity splicing between patches, producing a unified and detailed final mesh.
3.4 Training Strategy
Each mesh is partitioned into patches adaptively in training. Given an input mesh $\mathcal{M}$ with $\mathcal{N}_{p}$ vertices and $\mathcal{N}_{f}$ facets, the number of patches is set to $\mathcal{N}_{\text{seg}}=\frac{\mathcal{N}_{f}}{2000}\times\lambda_{\text{rand}}$, where $\lambda_{\text{rand}}$ is randomly sampled from [0.5, 2.5][0.5,\,2.5] to encourage diversity. The denominator ensures that each patch, after tokenization, yields a sequence length close to the window size ($9\mathrm{K}$) for efficient training.
We apply farthest point sampling (Moenning & Dodgson, 2003) to select $\mathcal{N}_{\text{seg}}$ points as cluster centers. Voronoi decomposition (Aurenhammer, 1991) partitions the mesh into patches based on these centers, with bisecting planes separating triangle regions. Clusters are ordered by breadth-first search, starting from the lowest center, to retrieve boundary information sequentially.
We also curate a subset of meshes with high-quality connected component annotations, using each component directly as a patch (with breadth-first ordering, as above). This subset supports tasks requiring more regular, consistent patch segmentation and enables training for semantic reasoning. We provide detailed training dataset analysis is in Appendix Sec. A.1.
4 Experiments
Table 1: Quantitative comparison on ShapeNet (Chang et al., 2015), Thingi10K (Zhou & Jacobson, 2016) and Objaverse (Deitke et al., 2023b) datasets. The best scores are emphasized in bold with underlining, while the second best scores are highlighted only in bold.
@@ 4 Experiments
Figure 8: Distribution of face count in our dataset.
4.1 Implementation
Our implementation builds upon the released $0.5B$ parameter DeepMesh (Zhao et al., 2025) model, which serves as the base for fine-tuning our approach. We introduce and progressively fuse new boundary conditions and global point cloud features into the architecture, connecting the GRU boundary encoder and global feature input using zero-initialized linear layers. Local point cloud features are mapped directly onto the original input cloud.
The curated dataset consists of 310K meshes, including approximately 90K with connected component information. The distribution of face count are visualized in Fig. 8. Please check Appendix Sec. A.1 for the data preprocess and more analysis. Training is conducted for seven days on a cluster of 32 NVIDIA H20 96GB GPUs, using a cosine learning rate scheduler that decays from $1\times 10^{-4}$ to $1\times 10^{-5}$. Token window sizes for truncated windows follow DeepMesh’s setting ($9K$, with a $50\%$ overlap). We employ KV-caching in both training and inference and adopt probabilistic sampling (temperature $0.5$) to ensure stable mesh generation.
4.2 Comparisons
To thoroughly assess the effectiveness of our proposed method, we perform comparative experiments with four publicly available state-of-the-art mesh generation methods: MeshAnythingV2 (Chen et al., 2024c), BPT (Weng et al., 2025), TreeMeshGPT (Lionar et al., 2025), and DeepMesh (Zhao et al., 2025). The comparison includes both quantitative measurements and qualitative visualization of results. We randomly select 100 samples from each of the ShapeNet (Chang et al., 2015), Thingi10K (Zhou & Jacobson, 2016), and Objaverse (Deitke et al., 2023b) datasets for all experiments.

### v2

# Rui Xu1  Tianyang Xue1  Qiujie Dong1  Le Wan2  Zhe Zhu2  Peng Li3

@@ Abstract
Scaling artist-designed meshes to high triangle numbers remains challenging for autoregressive generative models. Existing transformer-based methods suffer from long-sequence bottlenecks and limited quantization resolution, primarily due to the large number of tokens required and constrained quantization granularity. These issues prevent faithful reproduction of fine geometric details and structured density patterns. We introduce MeshMosaic, a novel local-to-global framework for artist mesh generation that scales to over 100K triangles—substantially surpassing prior methods, which typically handle only around 8K faces. MeshMosaic first segments shapes into patches, generating each patch autoregressively and leveraging shared boundary conditions to promote coherence, symmetry, and seamless connectivity between neighboring regions. This strategy enhances scalability to high-resolution meshes by quantizing patches individually, resulting in more symmetrical and organized mesh density and structure. Extensive experiments across multiple public datasets demonstrate that MeshMosaic significantly outperforms state-of-the-art methods in both geometric fidelity and user preference, supporting superior detail representation and practical mesh generation for real-world applications.
@@ 1 Introduction
Artist-designed triangular meshes are a cornerstone in film, gaming, AR/VR, and industrial design. High-quality artist meshes are central to computer graphics and 3D vision, distinguished by their stylized topology, directional flows, uneven triangle densities, sharp edges, and symmetry. Recent advances in 3D generation [TRELLIS, li2024era3d, liu2023syncdreamer, long2024wonder3d] and reconstruction [wang2023neural, 2dgs, GCNO, IPSR] highlight the limitations of classical meshing methods like Marching Cubes [MarchingCubes], which rely on uniform grids and produce redundant triangles, struggling with sharp features. Traditional meshing either yields uniform [cvt, xu2024cwf, PDT2025, dong2025neurcross, dong2025crossgen] or oversimplified [chen2023robust, qem] results; while anisotropic techniques [zhong2014anisotropic] better align to curvature, they still fall short in capturing the varying density and structure of artist meshes.
@@
The rise of large language models (LLMs) [zhao2023survey] has inspired GPT-like architectures for mesh generation, such as MeshGPT [Meshgpt] and its successors [chen2024meshxl, chen2024meshanythingv2, zhao2025deepmesh, hao2024meshtron, tang2024edgerunner]. Despite progress, these approaches struggle to scale up due to prohibitively long token sequences and limited quantization resolution, making it difficult to generate high-triangle meshes with fine detail. However, in practice, artist-designed meshes often require significantly higher resolutions to achieve the visual fidelity demanded in modern games and films. For example, production-quality character models or hero assets frequently contain upwards of 100K faces, far exceeding the capacities handled by current generative methods [zhao2025deepmesh, bpt, lionar2025treemeshgpt]. This substantial gap underscores the need for methods capable of generating high-triangle meshes that preserve the intricate details and structural coherence. Fig. 2 presents a qualitative comparison between our method and existing state-of-the-art approaches, including both academic and commercial models. Despite using a compact 0.5B parameter model, our results markedly outperform BPT [bpt] and DeepMesh [zhao2025deepmesh] in reconstruction and mesh quality. Very recently, Hunyuan3D [lei2025hunyuan3dstudio] released a commercial enhanced version built upon the BPT [bpt] framework, although the model size was not disclosed. Notably, our method attains comparable visual fidelity with only a small-scale 0.5B model and even surpasses it in terms of geometric completeness.
[…]
Figure 2: Comparison with existing state-of-the-art approaches, including both academic and commercial models. MeshMosaic achieved better quality with a smaller model size.
@@
Inspired by the compositional principles of classical mosaic art (Fig. 3), we propose MeshMosaic, a novel local-to-global framework for scalable artist mesh generation. Mosaic artworks achieve global complexity and coherence by assembling intricate local tiles; in a similar spirit, MeshMosaic constructs a complete mesh by stitching together multiple locally generated patches. Unlike previous methods that attempt to model the entire mesh sequence, our framework divides the mesh into semantically meaningful patches, each autoregressively generated from a full-size point cloud with full-resolution quantization.
Figure 3: Mosaic Art [mosaic].
To enable a compact yet faithful geometric representation, we employ shared boundary conditions and semantic segmentation, which address challenges related to boundary alignment and asymmetry. This patch-based strategy not only sidesteps the long-sequence bottleneck, but also effectively captures fine-grained geometric structures and global coherence, allowing for high-detail modeling within each patch while maintaining consistency across the entire mesh.
@@
Experiments on multiple datasets show that MeshMosaic establishes new milestones in geometric fidelity and detail, also strongly preferred in user studies for artistry. Our approach supports stable generation of high-resolution meshes with over 100K triangles (see Fig. 1) and faithfully reproduces fine detail via per-patch quantization.
@@ 2 Releated Works
Remarkable advances have been made in 3D shape generation, particularly with the adoption of signed distance field (SDF) representations, which offer notable accuracy and flexibility for modeling complex shapes. Despite such progress, these SDF-based methods often depend on the Marching Cubes algorithm [MarchingCubes] for mesh extraction, which can result in redundant triangles and consequently large file sizes—posing limitations for scalable deployment and real-time applications.
For instance, Wonder3D [long2024wonder3d] introduces a cross-domain diffusion framework for generating high-quality, multi-view textured 3D meshes from single images, achieving improved consistency and visual fidelity over previous approaches. CLAY [CLAY] expands the scope with a large-scale generative model that transforms text, images, and 3D-aware inputs into intricate geometry and material compositions, making robust 3D asset creation accessible to broad user bases. TRELLIS [TRELLIS] leverages structured occupancy fields to guide the formation of salient shape features, supporting high-precision modeling conditioned on text or image prompts. Hunyuan3D-2.5 [lai2025hunyuan3d] proposes a two-stage diffusion pipeline for crafting high-fidelity assets, combining powerful generative models with physically-based rendering for enhanced realism in both shape and texture. CraftsMan3D [li2024craftsman3d], evolves toward interactive 3D design by developing a native diffusion-based framework capable of producing meshes with regular topology and fine surface detail, while supporting user-driven refinements.
2.2 Artist Mesh Generation
The quest for artist-quality mesh generation has inspired a new wave of models that focus on efficient topology and expressive geometry. MeshGPT [Meshgpt] pioneers autoregressive mesh synthesis through sequence-based modeling, employing quantized latent embeddings and transformer architectures to predict efficient triangulation and structural patterns reminiscent of hand-crafted meshes. Building on this idea, MeshAnything [chen2024meshanything] and MeshAnythingV2 [chen2024meshanythingv2] offer advanced mesh generation using adjacent mesh tokenization, reducing token sequence lengths and enabling more complex, artist-grade meshes, with MeshAnythingV2 doubling the operational face limit.
MeshXL [chen2024meshxl] introduces the Neural Coordinate Field, which fuses explicit coordinate representation with implicit neural embeddings for more scalable, high-fidelity mesh modeling. EdgeRunner [tang2024edgerunner] addresses past limitations of autoregressive mesh approaches by presenting an improved tokenization algorithm and compressing variable-length meshes into fixed-size latent vectors, yielding more diverse, generalizable, and higher-quality outputs. Meshtron [hao2024meshtron] leverages a novel hourglass neural architecture with sliding window inference and robust sampling, achieving new levels of scalability and fidelity.
In addition, TreeMeshGPT [lionar2025treemeshgpt] introduces a tree sequencing method for triangle adjacency, dynamically growing mesh structures during autoregressive generation for improved training and mesh quality. iFlame [wang2025iflame] balances efficiency and generative power by combining linear and full attention within an hourglass framework, augmented by caching for fast inference and training. Nautilus [wang2025nautilus] explores locality-aware autoencoding by leveraging manifold mesh properties, novel tokenization, and dual-stream conditioning, significantly enhancing scalability and structural consistency.
Compression-oriented approaches such as Blocked and Patchified Tokenization (BPT) [bpt] further reduce token sequence length, allowing detailed mesh synthesis with more faces. Building on BPT, DeepMesh [zhao2025deepmesh] integrates reinforcement learning for human preference alignment, supporting the generation of intricately detailed meshes with precise topology. In addition to autoregressive-based approaches, methods such as PolyDiff [alliegro2023polydiff] and PDT [PDT2025] directly employ diffusion models to generate structured triangles or points from Gaussian noise.
@@ 2.2 Artist Mesh Generation
Figure 5: The pipeline of MeshMosaic. During inference, our method first applies PartField [liu2025partfield] to obtain semantic segmentation of the input shape. The input point cloud is then sampled according to the segmented patches and the original shape. Finally, our approach produces a clean, highly detailed mesh by assembling the generated patches.
@@ 2.3 Part-based Shape Generation
Part-based shape generation rests on the principle that decomposing objects into semantic parts furnishes rich priors for structure-aware reconstruction and controllable synthesis. Universal segmentation techniques have scaled part discovery across a wider range of data. Segment Any Mesh [tang2024segment] generalizes promptable segmentation to 3D meshes, supporting flexible and category-agnostic part extraction crucial for interactive and generative workflows. SAM3D [yang2023sam3d] adapts this paradigm to large-scale 3D scenes, enabling multi-granular, prompt-driven segmentation. By distilling knowledge from SAM’s multi-view segmentation results, SAMPart3D [yang2024sampart3d] further specializes in part segmentation for individual objects. More recently, PartField [liu2025partfield] represents shapes as continuous feature fields and trains a transformer-based feed-forward network with an ambiguity-agnostic contrastive loss, achieving efficient and high-quality open-world part segmentation. PartCrafter [lin2025partcrafter] jointly creates multiple semantically distinct parts from a single image, enabling end-to-end part-aware 3D mesh synthesis with global coherence and fine-grained detail.
Human modelers typically create models based on their understanding of component-based structures [lin2020modeling], and thus part-based generation is a problem of significant importance. Part123 [part123] illustrates this by reconstructing 3D shapes from single images while predicting semantic parts and their spatial arrangement. ComboStoc [ComboStoc2024] introduces combinatorial stochasticity into diffusion by jointly sampling discrete structural decisions (such as part templates or multiplicity) with continuous geometry. These segmentation frameworks underpin part-based shape generation by providing scalable, promptable part vocabularies and supporting interactive conditioning and evaluation at the part level.
3 Preliminaries
Recent works such as MeshGPT [Meshgpt], Meshtron [hao2024meshtron], BPT [bpt], and DeepMesh [zhao2025deepmesh] have pioneered the application of autoregressive transformers to the generation of artist meshes. These approaches convert unordered triangular meshes into ordered token sequences through specialized tokenization schemes. For example, DeepMesh [zhao2025deepmesh] selects high-degree vertices in ascending order along the vertical axis, encodes their first-order neighboring triangles as tokens, and iteratively proceeds to the next vertex according to its degree until the entire mesh is serialized into a token sequence (Fig. 4 left). Each point is subsequently represented within a quantized three-level block coordinate system, which enables compact and structured spatial encoding (Fig. 4 right).
Figure 4: DeepMesh Tokenizer.
This representation enables direct learning through autoregressive transformers. Meshtron [hao2024meshtron] also improves efficiency by employing hourglass transformer and introducing a sliding-window training strategy, in which long token sequences are segmented into fixed-length overlapping windows (50% overlap) to balance contextual coverage and computational cost.
@@ 4.2 Local-to-Global Mesh Generation
We use PartField [liu2025partfield] for semantic segmentation at inference time (see Fig. 5), which embeds semantic structure and produces well-aligned boundaries, often guided by curvature flow to enhance realism and make future edits easier.
Sorting patches.
Then, we will generate the whole mesh in a part-by-part manner, which requires us to determine a generation order. Given a pre-segmented shape, two patches are considered adjacent when they share at least one pair of neighboring triangles assigned to different segments.
@@
Thus, the patch generation is carried out in breadth-first search (BFS) order, beginning from the spatially lowest patch and then proceeding to adjacent patches. When the current patch has multiple neighbors, the one with the lowest coordinates is selected first to ensure a consistent and unique ordering. This strategy guarantees that each subsequent patch connects to at least one previously generated patch (except in the case of new connected components), enabling the propagation of critical boundary information to maintain structural symmetry and smoothness. Sequential generation with the autoregressive model ultimately yields the final mesh assembly. Fig. 6 shows a 2D illustration with eight patches, where the black line demonstrates the mesh surface. Then, for each patch, we adopt the following structure to generate the triangle meshes.
@@
Figure 7: The workflow of MeshMosaic for generating a single patch. Both global and local point cloud features are extracted by a locked Michelangelo [zhao2023michelangelo] encoder. For each patch, the nearest boundary mesh is identified, tokenized, and concatenated before the target mesh token sequence. The GRU network encodes boundary tokens, which are then combined with global and local features and fed into an autoregressive hourglass transformer for mesh generation.
@@ 4.3 Generating single Patch
When generating triangles for a specific patch, we will use the triangles from existing generated patches as boundary conditions. This essentially enables the smooth connection between different patches. Specifically, we introduce an efficient boundary conditioning mechanism: the token sequence from earlier patches is fed as context to subsequent patches. To avoid inefficiency and information dilution from excessively long token sequences, for each patch, we select only 512 spatially nearest triangles from prior patches (no segmentation produces more than this number of boundary triangles in our dataset). These triangles are tokenized by our tokenizer, passed through a Gate Recurrent Unit [GRU2014] (GRU) network, and the resulting embedding conditions the transformer network (see Fig. 7, blue arrows). Compared with simple pooling or fixed-length encoders, a GRU is particularly suitable for handling variable-length sequential inputs, as it can capture temporal dependencies and selectively retain boundary information over long contexts. This makes it well-suited for processing our boundary token sequences, whose lengths vary across patches depending on geometric complexity. In the upper half of Fig. 8, we illustrate an 2D example where a pentagon is divided into five patches. On the left, when generating the second patch, the nearest boundary information from the first patch is extracted to form a boundary token, which is then concatenated with the token sequence of the current patch. On the right, when the first four patches have been generated and the final patch is about to be produced, the nearest region from the existing patches is selected as the boundary condition for generation. Fig. 8 also shows a 3D example of the plane shape. The whole shape was segmented into three patches. Each patch gets the boundary information from the previous patches, following the bottom-to-top BFS order. For the very first patch, where no previous boundaries exist, we supply a placeholder token sequence consisting entirely of terminator tokens to the GRU network, establishing a neutral starting context for the generation process.
@@
Table 1: Quantitative comparison on ShapeNet [chang2015shapenet], Thingi10K [zhou2016thingi10k] and Objaverse [deitke2023objaverse] datasets. The best scores are emphasized in bold with underlining, while the second best scores are highlighted only in bold.
@@
Local-to-Global Conditioning.
While boundary conditioning enforces local continuity, we enhance global coherence via local-to-global point cloud features (middle of Fig. 7). During training and inference, our autoregressive model is conditioned on representations from both the current patch point clouds and the full shape point clouds. Both point clouds are encoded with a frozen Michelangelo [zhao2023michelangelo] encoder. The extracted global and local features are concatenated with GRU boundary features and provided to the transformer as final condition input.
Local Quantization.
Given both the boundary conditions and local-to-global information, we generate the triangles locally using a local quantization. Unlike previous approaches, such as DeepMesh [zhao2025deepmesh], which apply a uniform quantization of $512^{3}$ resolution to the entire mesh, our method independently scales each patch to $[0,1]$ and quantizes it at $512^{3}$ resolution. This local quantization approach enables a higher effective merged resolution, allowing for the preservation and recovery of richer geometric details. As shown in Fig. 5 and Fig. 7, our pipeline first segments the shape into patches, with each patch quantized independently to $512^{3}$ resolution and provided with 16,384 sampled points as input. In contrast to baseline methods, which quantize the full shape and use only a single set of 16,384 point cloud samples, our framework assembles meshes from individually quantized patches, each paired with its respective sampled points. This strategy offers the dual benefits of higher overall shape resolution and a greater abundance of conditional information for the network to leverage during generation.
Gluing local patches.
It should be noted that local quantization may introduce minor positional displacements for each patch, which can lead to discontinuities along patch seams if not properly addressed. To ensure seamless integration, we compute the displacement between the position of boundary condition faces referenced by the current patch and their corresponding original quantized positions in the previously assembled patches. The entire current patch is then translated according to this computed displacement, aligning it precisely with previously assembled patches. This compensatory adjustment guarantees smooth boundaries and continuity across the entire mesh, enabling high-fidelity splicing between patches, producing a unified and detailed final mesh. Because the boundary triangles between patches are replicated exactly, the gluing process is highly stable and computationally efficient.